OpenStudy (anonymous):

help me please? What is a quartic function with only the two real zeros given? x = -4 and x = -1 A. y = x^4 + 5x^3 + 5x^2 + 5x + 4 B. y = x^4 - 5x^3 - 5x^2 - 5x - 4 C. y = -x^4 + 5x^3 + 5x^2 + 5x + 4 D. y = x^4 + 5x^3 + 5x^2 + 5x - 5

OpenStudy (agent0smith):

Plug each of the x values into the functions, and check the whole thing is equal to zero... so start with x=-4 plugged into y = x^4 + 5x^3 + 5x^2 + 5x + 4 if it's equal to zero, and it's also zero when you plug in x=-1, that's your function.

OpenStudy (anonymous):

can you help me with other questions as well please

OpenStudy (agent0smith):

Maybe but i gotta go home very soon.

OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67) A. y = -x^3 + 4x^2 + 2x - 2 B. y = x^3 + 4x^2 + 2x - 2 C. y = x^4 + 4x^3 + 2x^2 - 2x D. y = x^3 - 4x^2 - 2x - 2

OpenStudy (anonymous):

Plug in the x, see if the y is the same as shown. If not, the point does not pass through

OpenStudy (anonymous):

All of the points must pass through the function for it to be true.

OpenStudy (anonymous):

how do i do that...?

OpenStudy (anonymous):

Take the x value for the point (-2,2) for the function A does it get the same y value?